ORDER: DOPAMINE 20MCG/KG/MIN AVAILABLE 800MG/500ML WEIGHT 165 LB HOW MANY ML/HR?
Answer #1

ORDER

DOPAMINE

NEEDED DOSE = 20 mcg / kg / min

AVAILABLE DOSE = 800 mg / 500 ml

WEIGHT OF THE PATIENT = 165 LB

TO CONVERT THE WEIGHT FROM LB TO KG   1 LB = 0.45359237 kg

165 lb X 0.45359237 kg   = 74.84 Kg ( weight of the patient in kg )

Amount of Drug Needed to Administer According to Weight

74.84 kg X 20 mcg / kg / min = 1496.8 mcg / min

To convert ,

1496.8 MCG to  MG   = 1496.8 / 1000 = 1.4968 mg / min.

in 1-hour total dose will be given = 1.4968 mg X 60 ( 1 HOUR) AS WE HAVE TO CALCULATE in ml/hr

The total dose needed in 1 hour = 89.808 Mg / Hr

to calculate the needed dose in ml

X = D / H * Q

WHERE

X= UNKNOWN DOSE IN ML

D = DESIRED DOSE ( 89.8 mg ) needed / hr

H= DOSE IN HAND ( 800 mg ) given

Q = VOLUME OF DILUTION ( 500 ml) given

X = D / H * Q

X= 89.8 Mg / 800 mg * 500 ml

X = 56.125 ML ( VOLUME NEEDED )

FLOW RATE = VOLUME (ml) / TIME ( HOUR)

= 56.125 ml / 1 Hr

= 56.125 ml / hr Amount of drug will be Administered in ONE hour.